A support for a roof truss is 2√5 feet long.
Which of the following is the same measurement?
A.5√2 feet
B.√10 feet
C.V20 feet
D.10 feet

: A support for a roof truss is 2√5 feet long.
Which of the following is the same measurement?
A.5√2 feet
B.√10 feet
C.V20 feet
D.10 feet


Answer by stanbon(18787)

A support for a roof truss is 2√5 feet long.
Which of the following is the same measurement?
---------------------------
2sqrt(5) = sqrt(4*5) = sqrt(20)
=======================
Cheers,
Stan H.